The rectilinear motion of a body with a mass of 2 kg is given by the law s= 2t +t^2.
What is the force acting on the body?
1
Expert's answer
2020-12-06T17:19:04-0500

Solution

Motion law for this body is given

"s=2t+t^2"

Now differentiate with respect to t

"v=\frac{ds}{dt}=2+2t"

Again differentiate with respect to t

Then acceleration

"a=\frac{dv}{dt}=\frac{d^2s}{dt^2}=2"

So force acting on the body.

"F=ma=2\times 2=4N"